OpenStudy (anonymous):

Hi guys, trying to figure out using Gaussian Elimination, but Can't seem to get into triangular form with second equation... the answer is (8,9,-1) but this is as far as I get in my work before getting stuck 3 -2 6 0 1 3 20 15 10 -11 -10 -9 swapping e2&e1 & -3(e1)+(e2) 1 3 20 15 0 -11 -54 -45 10 -11 -10 -9

hero (hero):

Add -10*R1 to R3

OpenStudy (anonymous):

I tried that but then it changes the value to r3's y value and I can't zero it out.? I just don't understand what Im doing wrong?

hero (hero):

Why are you afraid of changing R3's y value?

OpenStudy (anonymous):

Cause if I multiply -10 times r1 I zero out the 10 but then y becomes -41 and I can't figure out the constant to multiply r2 by to add to r3 to knock out the -41?

hero (hero):

You'll just have to multiply the third row by a number and then multiply the second row by a number so that when you combine them the 41 cancels to zero. Not that difficult

OpenStudy (anonymous):

Ok. Thank you. for the help I appreciate it.

hero (hero):

Is that all you needed?

OpenStudy (anonymous):

yea I was just giving it a shot. thank you

hero (hero):

Have you given up already? Or do you plan to finish this?

OpenStudy (anonymous):

no, I'm going to finish it. Im just going to see if I can find the multiple.

hero (hero):

Exactly. Find the least common multiple of -41 and -11

hero (hero):

And you can probably figure out what it is since both 11 and 41 are prime

OpenStudy (anonymous):

ahhhh.... ok.

OpenStudy (anonymous):

Uhhh, wait a second i'm still not getting it?.. I don't know what the least common multiple is, 1?

hero (hero):

That's the common factor, not the common multiple

OpenStudy (anonymous):

oh... 1/2

hero (hero):

The lcm of 11 and 41 is \(11 \times 41\)

OpenStudy (anonymous):

451 :) wow, I think i'm going to just go back to beginning algebra. ok, thank you. I guess I need more help than I think.

OpenStudy (anonymous):

thank you for your time.

hero (hero):

The gcf will always be less than or equal to either of two numbers of interest. The lcm will always be greater than or equal to either of the two numbers of interest

hero (hero):

\[\text{gcf} \le x \space\text{or}\space y \le \text{lcm}\]

OpenStudy (anonymous):

Awesome, I will copy and paste that into my notes, here. That's an easy way to remember it. Thanks. No matter how many times I think I can't figure it out, you guys are the best. Thank you.

hero (hero):

For example The gcf of 2 and 8 is 2 The lcm of 2 and 8 is 8

hero (hero):

Another example The gcf of 11 and 13 is 1 The lcm of 11 and 13 is 143

hero (hero):

Good luck with your studies

OpenStudy (anonymous):

Thank you.